OpenStudy (inkyvoyd):

Find the area: \(\Huge \int\limits_{0}^{\pi/8} \sqrt{\tan x} dx\) NOTE: Inky wants to inform you that this IS a troll question. You can choose to answer it, but it's definitely not worth inky's homework credit or inky's test grade.

OpenStudy (experimentx):

let tan(x) = u^2 dx = u/sqrt(1-u^2) du

OpenStudy (experimentx):

*2udu/sqrt(1-u^2)

OpenStudy (inkyvoyd):

the hard part is the partial fractions :3

OpenStudy (experimentx):

the indefintie integral appears to be in closed form. \[ \frac{2 \tan ^{-1}\left(\sqrt{2} \sqrt{\tan (x)}+1\right)-2 \tan ^{-1}\left(1-\sqrt{2} \sqrt{\tan (x)}\right)-\log \left(\tan (x)+\sqrt{2} \sqrt{\tan (x)}+1\right)+\log \left(\tan (x)-\sqrt{2} \sqrt{\tan (x)}+1\right)}{2 \sqrt{2}} \]

OpenStudy (inkyvoyd):

okay, you tried hard enough.

OpenStudy (experimentx):

haven't ... just woke up.

OpenStudy (experimentx):